OpenStudy (anonymous):

pls help... rationalize the denominator of 12/(3+sqrt{5}+2sqrt{2}

OpenStudy (anonymous):

\[12 \div ( 3 + \sqrt{5}) + 2\sqrt{2}\]

OpenStudy (anonymous):

Is that it?

OpenStudy (anonymous):

Is that the question ?

OpenStudy (anonymous):

yes...but no brckt aftr root5 ...afrt 2 root 2

OpenStudy (anonymous):

can u pls help mee...plssss...

OpenStudy (anonymous):

yes that makes no difference Ignore the numerator for now 3 + \[\sqrt{5} + 2\sqrt{2}\]

OpenStudy (anonymous):

take it as 3 + ( sqrt 5 + 2 sqrt 2 ) Miltiply this by ( sqrt 5 - 2 sqrt 2 / sqrt 5 - 2 sqrt 2)

OpenStudy (anonymous):

should we take 3 ??

OpenStudy (anonymous):

i got\[12\sqrt{5}-24\sqrt{2}/3\sqrt{5}-6\sqrt{2}-3\] is that correct??

OpenStudy (amistre64):

we really should get the sqrts haveing the same radicals and innards.... if possible right?

OpenStudy (amistre64):

or maybe....

OpenStudy (anonymous):

thnx evry1 !!